OpenStudy (anonymous):

I just need an explanation PLEASE!! Tan(77)=x+y/200. I'm not sure how the teacher got the answer but he said the answer was x+y=866.3.

OpenStudy (loser66):

\(tan (77) =\dfrac{x+y}{200}\\200tan(77) = x+y\) , hit some keys in your calculator to find tan(77) , you get what your prof got

OpenStudy (anonymous):

Okay, thank you!